OpenStudy (anonymous):

(√5 + 3√8)(√5 - 3√8) I have an ans. -67 plse check

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

LagrangeSon678 plse check Pls check: use rational exponents to write single radical expression: ^4√8 * ^2√7 I have four ans to choose from: I came up with ^4√392

OpenStudy (anonymous):

What are the ^ in front of the 4 and the 2 suppose to be?

OpenStudy (anonymous):

I'll use the equation editior

OpenStudy (anonymous):

\[\sqrt[4]{8} * \sqrt[2]{7}=\]

OpenStudy (anonymous):

the symbol at the end is a equal sign

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what other answers are there? cause i got a different one from yours

OpenStudy (anonymous):

my four choices of ans. to choose from is:\[\sqrt[4]{175616}\]\[\sqrt[8]{56}\]\[\sqrt[4]{392}\]\[1/392^{4}\]

OpenStudy (anonymous):

i got : 2^3/4 *sqrt 7

OpenStudy (anonymous):

are you sure you wrote the problem correctly?

OpenStudy (anonymous):

double check

myininaya (myininaya):

its the third choice

myininaya (myininaya):

\[8^\frac{1}{4}*7^\frac{1}{2}=8^\frac{1}{4}*7^\frac{2}{4}=(8*7^{2})^\frac{1}{4}=392^\frac{1}{4}\]